# Investigating the Robustness of Extreme Precipitation Super-Resolution Across Climates

### v2

@@ 5 Results
We first investigate on present-day data the performance of our super-resolution models from $13.2$ km to $2.2$ km trained on present-day data. We then study how the performance of our super-resolution models varies when decreasing the resolution of the input data to which they are trained ($26.4$ and $52.8$ km, instead of $13.2$), and explain the observed behavior using the tools introduced in Section 3.2. Finally, we investigate, using the methodology developed in Section 3.3, the robustness of our super-resolution models in a warmer climate.
@@ 5.1 Performance in the Reference Climate
Table 1: Comparison of VGAM and VGLM performance across training and test sets using mean (across grid points) Cramér–von Mises errors. The VGAM consistently outperforms the VGLM in the reference climate.
@@ 5.2 Generalization across Super-Resolution Factors
Over the Swiss Alps, the 52.8 km model using GEV features underperforms compared to an elevation-only model, reflecting the loss of predictive content in coarse precipitation fields. At such resolutions, alternative downscaling methods—such as perfect-prognosis approaches that incorporate temperature or humidity—or deep learning architectures exploiting broader spatial contexts may improve performance.
@@
To better understand the observed performance drop, we apply the methdology described in Section 3.2 to the the 2.2 km precipitation and elevation fields. More precisely, we investigate the behavior of the spatial correlation and autocorrelation functions defined in (7) and (10). As illustrated in Figure 7a, the spatial correlation of precipitation decreases approximately exponentially with distance, whereas the spatial cross-correlation between precipitation and elevation remains nearly constant. These behaviors can be modeled using the following functional forms, as shown in Figure 7b:
@@ 5.3 Understanding Robustness to Climate Change
for the detailed derivation of these terms, see Sections 2.2 and 2.3 of the SM.
By incorporating this decomposition into (3.3), we obtain the decomposed form of the pointwise robustness gap $G$ in the GEV case:
@@
Panel (a) of Figure 9 shows that the pointwise robustness gaps increase markedly at higher quantile levels (above the 90th percentile), highlighting a clear limitation in the model’s ability to generalize under climate change. Panel (b) reveals that this degradation is largely driven by increasing contributions from the scale and shape parameters. This underscores the importance of accounting for potential shifts in these components. In particular, the shape parameter was kept fixed across space in our VGAM setup, which limits its capacity to capture changes in the tails of the distribution. However, this finding must be interpreted with caution: the extremes being analyzed are already the upper tail of an extreme value distribution, and the underlying sample size remains limited. This also explains the slight increase observed in the residual term.
Interestingly, for lower quantile levels (below the 30th percentile), the VGLM generalizes better, as indicated by consistently smaller pointwise robustness gaps. Panel (b) reveals that the location parameter term $T_{\mu}$ plays a dominant role. For the VGAM, its contribution rises earlier and more sharply than for the VGLM, as shown by the gap between the solid and dashed blue lines. In such cases, the location parameter accounts for nearly all of the robustness behavior, while the contributions from scale and shape remain close to zero.
@@
6 Conclusion
In this work, we introduced a framework for super-resolving distributions within interpretable statistical models, focusing on risk-relevant quantities such as high quantiles (return levels). This approach is particularly valuable for impact modeling, as it enables the estimation of risk measures at resolutions suited to downstream applications. We developed a method to identify the resolution scale at which super-resolution begins to fail and proposed the novel concept of a robustness gap, which we analyzed in the context of quantile estimation using the pinball loss. Applied in a pseudo-reality experiment over Switzerland, our framework showed that the quantile-wise robustness gap is an effective diagnostic for evaluating how well models trained on present-day data generalize to warmer climates. Leveraging a GEV distribution parameterized via VGAMs, we could pinpoint which model components–parameters and splines–contribute to robustness failures, offering insights into model limitations and guiding future improvements. Overall, the proposed methodology enables a tractable decomposition of spatial generalization and robustness across climates.
Several directions could further improve this framework. First, our use of mean pooling to approximate coarse-resolution fields is a first-order simplification; future work could explore alternative filters (e.g., Gaussian) or incorporate temporal and physical biases through explicit bias correction. Second, we assumed a constant shape parameter for the GEV distribution, which may not reflect how extremes respond to climate shifts. Using the $r$-largest values approach could stabilize tail estimates and support state-dependent shape parameter modeling, helping address an open question: which covariates most strongly influence the shape parameter $\xi$? Third, since the original pseudo-global warming simulation used here (Hentgen et al., 2019), the pseudo-global warming method has been refined (Brogli et al., 2023; Heim et al., 2023), and future work should update the analysis using the latest available simulations. Finally, tailored versions of the GEV, such as the blended GEV (Vandeskog et al., 2022), may improve the stability and realism of tail estimates.
Overall, our results suggest that combining probabilistic super-resolution with quantifiable generalization/robustness diagnostics provides a principled framework for modeling extremes under distributional shifts, with potential relevance to other fields that study rare events in non-stationary systems.
@@ Acknowledgments
We thank Marine Berthier and Leo Micollet for preliminary work that laid useful foundations for this study, Linda Mhalla and Shivanshi Asthana for feedback that improved our manuscript, and Laureline Hentgen, Nikolina Ban and Christoph Schär, for performing the kilomter scale PGW simulations. These simulations were conducted on the Piz Daint supercomputer of the Swiss National Supercomputing Centre (CSCS) under project ID pr144. EK would like to thank the Expertise Center for Climate Extremes (ECCE) at UNIL for financial support. TB acknowledges partial support from the Swiss National Science Foundation (SNSF) under Grant No. 10001754 (“RobustSR” project).

### v3

@@ 4.3 Selecting Features and Hyperparameters
Finally, we did not include temporal trend terms in the model because they were not statistically significant (see Section D in the SM for details).
@@ 5 Results
We first evaluate models trained in the present (CTRL) climate that super-resolve $13.2$ km inputs to $2.2$ km targets. We then assess how performance changes as inputs are coarsened to $26.4$ and $52.8$ km (super-resolution factors of $12\times$ and $24\times$) instead of $13.2$ km, and explain the observed behavior using the tools introduced in Section 3.2. Finally, we investigate, using the methodology developed in Section 3.3, the robustness of our super-resolution models in a warmer climate.
@@ 5.1 Performance in the Reference Climate
Table 1: Comparison of VGAM and VGLM performance based on mean Cramér–von Mises errors across grid points. The first three columns show results for the training, validation, and test subsets within the climate used for model training (present or future). The last column reports performance over the entire domain in the climate not used for training, assessing spatial generalization across climates. VGAM consistently outperforms VGLM in the present (reference) climate.
@@ 5.2 Generalization across Super-Resolution Factors
Over the Swiss Alps, the 52.8 km model using GEV features underperforms compared to an elevation-only model, reflecting the loss of predictive content in coarse precipitation fields. At such resolutions, alternative downscaling methods—such as perfect-prognosis approaches incorporating temperature or humidity, or deep learning architectures leveraging broader spatial contexts—may enhance performance.
@@
To better understand the observed performance drop, we apply the methodology described in Section 3.2 to the the 2.2 km precipitation and elevation fields. More precisely, we investigate the behavior of the spatial correlation and autocorrelation functions defined in (7) and (10). As illustrated in Figure 6a, the spatial correlation of precipitation decreases approximately exponentially with distance, whereas the spatial cross-correlation between precipitation and elevation remains nearly constant. These behaviors can be modeled using the following functional forms, as shown in Figure 6b:
@@ 5.3 Understanding Robustness to Climate Change
for the detailed derivation of these terms, see Sections C.2 and C.3 of the SM.
By incorporating this decomposition into (3.3.2), we obtain the decomposed form of the pointwise robustness gap $G$ in the GEV case:
@@
Panel (a) of Figure 8 shows very large pointwise robustness gaps at high quantile levels (above the 90th percentile), revealing clear limitations in the model’s ability to generalize under climate change. Panel (b) indicates that this degradation is mainly driven by the scale and shape parameters, highlighting the importance of accounting for changes in dispersion and tail behavior. In particular, the assumption of a spatially constant shape parameter in the VGAM strongly limits the model’s capacity to capture modifications in the distribution tails under future conditions.
This interpretation must, however, be nuanced. In the normalized robustness gap (Figure 9), the sharp increase at high quantiles is clearly attenuated, indicating that the large absolute gaps occur in a regime where the model trained on future already exhibits high losses. This reflects the intrinsic difficulty of predicting very rare extremes rather than solely a lack of model transferability.
At lower quantiles, Figure 8(a) shows better generalization of the VGLM, with consistently smaller absolute pointwise robustness gaps. In contrast, Figure 9(a) reveals relatively large normalized gaps at the 5th and 10th percentiles, mainly due to the very low losses of the future model in this regime. Across both figures, Panel (b) consistently shows that the location parameter term $T_{\mu}$ dominates the robustness gap, while scale and shape contributions remain negligible, underscoring the central role of the location parameter in governing robustness at low quantiles.
@@
Figure 9: Same as Figure 8, with all values—pointwise robustness gaps in (a) and contributions in (b)—normalized by $\ell_{\alpha}\left(q_{\text{F},\alpha},q^{\text{F}}_{\text{F},\alpha}\right)$. The normalization of the contributions in (31) is indicated by adding a tilde symbol.
6 Summary and Conclusions
6.1 Summary
As illustrated in Figure 10, we introduced a framework for super-resolving distributions within interpretable statistical models, focusing on risk-relevant quantities such as high quantiles (return levels). This approach is particularly valuable for impact modeling, as it enables the estimation of risk measures at resolutions suited to downstream applications. We developed a method to identify the resolution scale at which super-resolution begins to fail and proposed the novel concept of a robustness gap, which we analyzed in the context of quantile estimation using the pinball loss. Applied in a pseudo-reality experiment over Switzerland, our framework showed that the quantile-wise robustness gap is an effective diagnostic for evaluating how well models trained on present-day data generalize to warmer climates. Leveraging a GEV distribution parameterized via VGAMs, we could pinpoint which model components–parameters and splines–contribute to robustness failures, offering insights into model limitations and guiding future improvements. Overall, the proposed methodology enables a tractable decomposition of spatial generalization and robustness across climates. Our framework is distribution-agnostic and extends beyond precipitation to variables with tractable distributions (e.g., temperature extrema, wind speeds/gusts, and river discharge). By contrast, choices such as using a GEV distribution for block-maxima, elevation-based covariates, and mean pooling for coarsening are precipitation/region-specific and may be adapted per variable and application.
6.2 Outlook and Conclusions
Several directions could further improve this framework. First, our use of mean pooling to approximate coarse-resolution fields is a first-order simplification; future work could explore alternative filters (e.g., Gaussian) or incorporate temporal and physical biases through explicit bias correction. Second, we assumed a constant shape parameter for the GEV distribution, which may not reflect how extremes respond to climate shifts. Using the $r$-largest values approach could stabilize tail estimates and support state-dependent shape parameter modeling, helping address an open question: which covariates most strongly influence the shape parameter $\xi$? Third, since the original pseudo-global warming simulation used here (Hentgen2019data), the pseudo-global warming method has been refined (Brogli_2023; Heim_2023), and future work should update the analysis using the latest available simulations. Finally, tailored versions of the GEV, such as the blended GEV (vandeskog2022modelling), may improve the stability and realism of tail estimates.
Overall, our results suggest that combining probabilistic super-resolution with quantifiable generalization/robustness diagnostics provides a principled framework for modeling extremes under distributional transformations, with potential relevance to other fields that study rare events in non-stationary systems.
@@ Acknowledgments
We thank Marine Berthier and Leo Micollet for preliminary work that laid useful foundations for this study, Linda Mhalla and Shivanshi Asthana for feedback that improved our manuscript, and Laureline Hentgen, Nikolina Ban and Christoph Schär, for performing the kilomter scale PGW simulations. These simulations were conducted on the Piz Daint supercomputer of the Swiss National Supercomputing Centre (CSCS) under project ID pr144. TB acknowledges support from the Swiss National Science Foundation (SNSF) under Grant No. 10001754 (“RobustSR” project). EK would like to thank the Expertise Center for Climate Extremes (ECCE) at UNIL for financial support.